OpenStudy (jenniferjuice):

Identify the correct slope and y intercept of the equation 2x + 4y = 12.

OpenStudy (anonymous):

kellin quin!

OpenStudy (anonymous):

YOu will need to change the equation to slope intercept form

OpenStudy (anonymous):

i was talking bout the pic lol

OpenStudy (anonymous):

I know, I love him. one of the band his is in song saved my life

OpenStudy (ja1):

Start by getting y by itself, what slope do you get?

OpenStudy (ja1):

\[2x + 4y = 12\] how do you get Y by itself? You take of the x variable and so you have to subrtact it form both sides.

OpenStudy (ja1):

Can you do that @jenniferjuice

OpenStudy (ja1):

JA1 feels the love V.V

OpenStudy (anonymous):

y=3-1/2x O.o?

OpenStudy (anonymous):

I haven't done it but what were your steps?

OpenStudy (texaschic101):

2x + 4y = 12......you want it in y = mx + b form subtract 2x from both sides 2x - 2x + 4y = 12 - 2x (combine like terms) 4y = -2x + 12 (now divide by 4 to get y by itself) 4y/4 = (-2x + 12)/4 y = -2/4x + 12/4 y = -1/2x + 3 Now your slope is -1/2 and your y intercept is 3

OpenStudy (anonymous):

correct, good job :)

OpenStudy (jenniferjuice):

@melody16 and @abashir you both have great taste in music